OpenStudy (anonymous):

find the value of 2.4 power by 2

OpenStudy (saifoo.khan):

\[\Large 2.4^ 2 = 2.4 *2.4= 4.8\]

OpenStudy (anonymous):

2.4\[2.4^{2}= 4.8\]

OpenStudy (anonymous):

Because it is to the second power. You have to multiply the number twice.

OpenStudy (anonymous):

2.4 x 2.4=5.76 i thought

OpenStudy (anonymous):

jackjones , you are right

OpenStudy (saifoo.khan):

OHO! he's correct. sorrie.

OpenStudy (anonymous):

plus, times, whatever